OpenStudy (anonymous):

What is the solution of 6(k - 2) = 3k - 6? -6 2 -4/3 2/3

OpenStudy (anonymous):

6k-12=3k-6 3k=6 k=3 bt why it is not in choices?:P

OpenStudy (anonymous):

mmm :/ does are the choices I have! is there any other chance it could be one of those choices?

OpenStudy (anonymous):

those*

OpenStudy (anonymous):

ok i will try just wait and watch :)

OpenStudy (anonymous):

k-2=k/2-1 k-k/2=1 k/2=1 k=2

OpenStudy (anonymous):

just divide by 6 to all eqn

OpenStudy (anonymous):

thanks! :)